# Video-Language Alignment Pre-training via Spatio–Temporal Graph Transformer

### v1

@@ 3.4 Self-Similarity Alignment Loss
To explore the similarity relationships in cross video-text pairs for further enhancing video-text alignment, we propose a self-similarity alignment loss with the initial optimization by contrastive learning. In initial stage, we employ contrastive learning loss ($\mathcal{L}_{vtc}$) to guide the model’s optimization, focusing on the point-to-point relationships between annotated video-text pairs. The $\mathcal{L}_{vtc}$ helps the model understand the basic video-language alignment, thereby facilitating convergence. The form of $\mathcal{L}_{vtc}$ is as follows:
@@
In self-similarity alignment stage, we adopt the self-similarity alignment loss ($\mathcal{L}_{ssa}$) to further refine video-language alignment by leveraging the inherent self-similarity present in the video and text data. The formula of the proposed $\mathcal{L}_{ssa}$ is as follows:
@@
$\mathcal{L}_{ssa}=-\frac{1}{2B}{\sum_{i=1}^{B}\sum_{j=1}^{B}\frac{e^{{\gamma}\cdot S_{{}_{v\cdot t}}^{(i,j)}}}{\sum_{k=1}^{B}e^{{\gamma}\cdot S_{{}_{v\cdot t}}^{(i,k)}}}(log(P_{{}_{v2t}}^{(i,j)}})+log(P_{{}_{t2v}}^{(i,j)})),$ (20)
where $S(v,v$ and $S(t,t)$ are the similarity between video-video and text-text, respectively. Unlike $\mathcal{L}_{vtc}$, the $\mathcal{L}_{ssa}$ focuses on distribution-to-distribution relationships by leveraging the inherent self-similarity within the video and text data, as shown in Fig. 5. With the $\mathcal{L}_{ssa}$, the alignment accuracy and overall performance are both improved. The hyper-parameter $\gamma$ must be greater than zero.
In addition to video-text contrastive loss $\mathcal{L}_{vtc}$ and the self-similarity alignment loss $\mathcal{L}_{ssa}$, we also use the vision-text matching loss $\mathcal{L}_{vtm}$ and vision-grounded text generation loss $\mathcal{L}_{vtg}$, following the approach BLIP-2 [31]. Consequently, the total loss of our method can be expressed as follows:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{vtc}+(1-\alpha)\cdot\mathcal{L}_{ssa}+\mathcal{L}_{vtm}+\mathcal{L}_{vtg}$ (21)
where parameter $\alpha$ is set to one in initial training stage and zero in self-similarity alignment training stage.
@@ 4.1 Datasets and Downstream Tasks
DiDeMo [58] contains 10k videos with 40k text descriptions. We evaluate video-paragraph retrieval following [30], where all descriptions for a video are concatenated into a single query.
LSMDC [59] consists of 118,081 video clips with one caption corresponding to each clip. Evaluation is conducted on a test set of 1,000 videos disjoint from the train and validation sets.
MSVD [60] contains 1,970 videos. Train, validation and, test splits contain 1,200, 100, and 670 videos, respectively. Each video has approximately 40 associated sentences in English.
(b) Video Question Answering
MSVD-QA [30] is built upon videos and text descriptions from MSVD. The MSVD-QA dataset has in total 1,970 videos and 50k question answer pairs, with 2,423 answer candidates.
@@ 4.3 Ablation Study
We conduct ablation experiments on MSRVTT to evaluate the effects of hyperparameters $\hbar$ and $\gamma$ in $\mathcal{L}_{ssa}$, as well as the STGT module and $\mathcal{L}_{ssa}$. These experiments focus on zero-shot text-to-video retrieval, using only Webvid-10M for pre-training.
@@
Temperature Coefficient $\gamma$. After the initial training with contrastive learning loss, we use the self-similarity alignment loss ($\mathcal{L}_{ssa}$) to further optimize the model. The choice of $\gamma$ significantly influences the performance of zero-shot text-to-video retrieval on MSRVTT. To explore this, we conduct experiments with integer values of $\gamma$ ranging from 1 to 12. As shown in Fig. 6 (b), the performance (R-Mean) initially increases and then decreases as $\gamma$ increases. When $\gamma$ is less than 3, $\mathcal{L}_{ssa}$ does not provide any significant gains. However, when $\gamma$ is between 3 and 10, $\mathcal{L}_{ssa}$ contributes to a notable improvement in performance. Once $\gamma$ exceeds 10, the performance improvement becomes negligible because the $\mathcal{L}_{ssa}$ and contrast loss ($\mathcal{L}_{vtc}$) become almost equivalent in this case.
TABLE III: Ablation study of STGT and $\mathcal{L}_{ssa}$.
Methods STGT $\mathcal{L}_{ssa}$ R@1$\uparrow$ R@5$\uparrow$ R@10$\uparrow$ R-$\uparrow$
@@
STGT and $\mathcal{L}_{ssa}$. As listed in Tab. II, the results demonstrate that the spatio-temporal graph ($\mathcal{A}$) brings a significant performance improvement of 3.29% in R-Mean. Additionally, the weight information ($\mathcal{W}s$) brings a slight improvement of 0.71%. Overall, the STGT achieves an impressive 4.1% improvement in R-Mean. Furthermore, self-similarity alignment loss ($\mathcal{L}_{ssa}$) brings a considerable 1.6% improvement in R-Mean, as well as comprehensive improvement across the board for R@1, R@5, and R@10 metrics.
@@ 4.4 Comparing to State-of-the-art
Video question answering. Tab. V shows the comparative results of our method and existing state-of-the-art (SOTA) methods on two video QA datasets. The models STGT+OPT2.7B and STGT+Vicuna7B outperform other methods on both datasets, demonstrating our method superior performance. Notably, when compared with the current SOTA method, Clover [13], our STGT+OPT2.7B model achieves significant improvements of 7.1% and 7.9% on the MSRVTT-QA and MSVD-QA, respectively. Similarly, STGT+Vicuna7B shows impressive improvements of 14.7% and 17.3% on the MSRVTT-QA and MSVD-QA, respectively. These experimental results further highlight the superiority and effectiveness of our method.
@@
Model parameter and efficiency analysis. We perform comparative experiments for our method and ALPRO [30] on 8 A100 GPUs. As listed in Tab. VI, despite a slight increase in parameters by 16M, our method nearly doubles ALPRO’s efficiency due to sparse design of spatio-temporal graphs transformer, showcasing its superior performance. In other words, our STGT module is straightforward and does not significantly increase complexity.
@@ 4.5 Visual Analysis of Experimental Results
Retrieval Comparison. In Fig. 12 and Fig. 13, we present a visual comparison of the retrieval results between our spatio-temporal graph transformer (STGT) and the ’Baseline’ model. In Fig. 12, the ’Baseline’ model, influenced by the visual similarity of videos, identifies three origami videos (Rank 1, Rank 3, and Rank 4) and two handicraft videos (Rank 2 and Rank 5). However, the correct video is only placed at Rank 4. In contrast, our model, enhanced with STGT and self-similarity alignment loss, accurately identifies four origami videos (Rank 1 4) and one handicraft video (Rank 5), with the correct video securing the top rank. This clearly illustrates the superior accuracy of our model.
@@
Effect of hyperparameter $\hbar$ on Spatio-Temporal Graph and Local Attention. Fig. 14 provide a visual representation of the attention distribution for reference tokens under various hyperparameter ($\hbar$) settings. In Fig. 14, the model’s capacity to differentiate between ‘cat’ and ‘dog’ reference tokens is displayed across a range of $\hbar$ values. At lower $\hbar$ values, the attention is more dispersed, indicating a wider, albeit less precise, focus on the relevant subjects within the image. As $\hbar$ increases, the attention distribution becomes more focused, suggesting a refined concentration that enhances the model’s ability to accurately distinguish between the cat and dog within the scene. However, it’s important to note that if $\hbar$ is excessively large, the reference local token may not effectively capture crucial local information between adjacent frames. This could result in under utilization of the temporal information between video frames, potentially leading to a decrease in the model’s feature table modeling ability. These figures collectively highlight the delicate balance required when setting the value of $\hbar$ to optimize the model’s performance in identifying and focusing on specific subjects within complex visual scenes.
@@ 5 Conclusion
This paper introduces a novel the spatio-temporal graph transformer (STGT) module for video-language alignment. By combining the strengths of both graph and transformer, the STGT effectively learns the spatial and temporal features of videos, utilizing the spatio-temporal contexts provided by the spatio-temporal graph. Additionally, a novel self-similarity alignment loss is used to explore the similarity relationships in cross video-text pairs for further enhancing the robustness of video-text alignment. Experimental results on challenging downstream tasks demonstrate that our method surpasses existing state-of-the-art methods, achieving the impressive performance in video-text retrieval and video question answering tasks.

### v2

@@ 3.4 Cross-similarity Alignment Loss
To explore the similarity relationships in cross video-text pairs for further enhancing video-text alignment, we propose a cross-similarity alignment loss with the the initial optimization by contrastive learning. In initial stage, we employ contrastive learning loss ($\mathcal{L}_{vtc}$) to guide the model’s optimization, focusing on the point-to-point relationships between annotated video-text pairs. The $\mathcal{L}_{vtc}$ helps the model understand the basic video-language alignment, thereby facilitating convergence. The form of $\mathcal{L}_{vtc}$ is as follows:
@@
In cross-similarity alignment stage, we adopt the cross-similarity alignment loss ($\mathcal{L}_{csal}$) to further refine video-language alignment by leveraging the inherent cross-similarity present in the video and text data. The formula of the proposed $\mathcal{L}_{csal}$ is as follows:
@@
$\mathcal{L}_{csal}=-\frac{1}{2B}{\sum_{i=1}^{B}\sum_{j=1}^{B}\frac{e^{{\gamma}\cdot S_{{}_{v\cdot t}}^{(i,j)}}}{\sum_{k=1}^{B}e^{{\gamma}\cdot S_{{}_{v\cdot t}}^{(i,k)}}}(log(P_{{}_{v2t}}^{(i,j)}})+log(P_{{}_{t2v}}^{(i,j)})),$ (20)
where $S(v,v$ and $S(t,t)$ are the similarity between video-video and text-text, respectively. Unlike $\mathcal{L}_{vtc}$, the $\mathcal{L}_{csal}$ focuses on distribution-to-distribution relationships by leveraging the inherent cross-similarity within the video and text data, as shown in Fig. 5. With the $\mathcal{L}_{csal}$, the alignment accuracy and overall performance are both improved. The hyper-parameter $\gamma$ must be greater than zero.
In addition to video-text contrastive loss $\mathcal{L}_{vtc}$ and the cross-similarity alignment loss $\mathcal{L}_{csal}$, we also use the vision-text matching loss $\mathcal{L}_{vtm}$ and vision-grounded text generation loss $\mathcal{L}_{vtg}$, following BLIP-2 [33]. Consequently, the total loss of our method can be expressed as follows:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{vtc}+(1-\alpha)\cdot\mathcal{L}_{csal}+\mathcal{L}_{vtm}+\mathcal{L}_{vtg}$ (21)
where parameter $\alpha$ is set to one in the initial training stage and zero in cross-similarity alignment training stage.
@@ 4.1 Datasets and Downstream Tasks
DiDeMo [61] contains 10k videos with 40k text descriptions. We evaluate video-paragraph retrieval following [32], where all text for a video are combined into a single query.
LSMDC [62] consists of 118,081 video clips with one caption corresponding to each clip. Evaluation is conducted on a test set of 1,000 videos.
MSVD [63] contains 1,970 videos. Train, validation, test splits contain 1,200, 100, and 670 videos, respectively. Each video has about 40 associated sentences in English.
MSVD-QA [32] is built upon videos and text from MSVD. The MSVD-QA has a total 1,970 videos and 50k question answer pairs, with 2,423 answer candidates.
@@ 4.3 Ablation Study
We conduct ablation experiments on MSRVTT to evaluate the effects of $\hbar$ and $\gamma$ in $\mathcal{L}_{csal}$, as well as the STGT module and $\mathcal{L}_{csal}$. These experiments focus on zero-shot text-to-video retrieval, only Webvid-10M for pre-training.
@@
Temperature Coefficient $\gamma$. After the initial training with contrastive learning loss, we use the cross-similarity alignment loss ($\mathcal{L}_{csal}$) to further optimize the model. The choice of $\gamma$ significantly influences the performance of zero-shot text-to-video retrieval on MSRVTT. Hence, we conduct experiments with integer values of $\gamma$ ranging from 1 to 12. As shown in Fig. 6 (b), the performance (R-Mean) initially increases and then decreases as $\gamma$ increases. When $\gamma$ is less than 3, $\mathcal{L}_{csal}$ provides minimal gains. However, when $\gamma$ is between 3 and 10, $\mathcal{L}_{csal}$ contributes to a notable improvement in performance. Once $\gamma$ exceeds 10, the performance improvement becomes negligible because the $\mathcal{L}_{csal}$ and contrast loss ($\mathcal{L}_{vtc}$) become almost equivalent in this case.
TABLE III: Ablation study of STGT and $\mathcal{L}_{csal}$.
Methods STGT $\mathcal{L}_{csal}$ R@1$\uparrow$ R@5$\uparrow$ R@10$\uparrow$ R-$\uparrow$
@@
STGT and $\mathcal{L}_{csal}$. As listed in Tab. II, the results demonstrate that the spatio-temporal graph ($\mathcal{A}$) brings a significant performance improvement of 3.29% in R-Mean. Additionally, the weight information ($\mathcal{W}s$) brings a slight improvement of 0.71%. Overall, the STGT achieves an impressive 4.1% improvement in R-Mean. Furthermore, cross-similarity alignment loss ($\mathcal{L}_{csal}$) brings a considerable 1.6% improvement in R-Mean, as well as comprehensive improvement across the board for R@1, R@5, and R@10 metrics.
@@ 4.4 Comparing to State-of-the-art
Video question answering. Tab. V shows the comparative results of our method and existing state-of-the-art (SOTA) methods on two video QA datasets. The models STGT+OPT2.7B and STGT+Vicuna7B outperform other methods on both datasets, demonstrating our method’s superior performance. Notably, when compared with the current SOTA method, Clover [12], our STGT+OPT2.7B model achieves significant improvements of 7.1% and 7.9% on the MSRVTT-QA and MSVD-QA, respectively. Similarly, STGT+Vicuna7B shows impressive improvements of 14.7% and 17.3% on the MSRVTT-QA and MSVD-QA, respectively. These experimental results further highlight the superiority and effectiveness of our method.
@@
Model parameter and efficiency analysis. We perform comparative experiments for our method and ALPRO [32] on 8 A100 GPUs. As listed in Tab. VI, despite a slight increase in parameters by 16M, our method nearly doubles ALPRO’s efficiency due to the sparse design of spatio-temporal graphs transformer, showcasing its superior performance. In other words, our STGT module is straightforward and does not significantly increase complexity.
@@ 4.5 Visual Analysis of Experimental Results
Retrieval Comparison. In Fig. 12 and Fig. 13, we present a visual comparison of the retrieval results between our spatio-temporal graph transformer (STGT) and the ’Baseline’ model. In Fig. 12, the ’Baseline’ model, influenced by the visual similarity of videos, identifies three origami videos (Rank 1, Rank 3, and Rank 4) and two handicraft videos (Rank 2 and Rank 5). However, the correct video is only placed at Rank 4. In contrast, our model, enhanced with STGT and cross-similarity alignment loss, accurately identifies four origami videos (Rank 1 4) and one handicraft video (Rank 5), with the correct video securing the top rank. This clearly illustrates the superior accuracy of our model.
@@
Effect of hyperparameter $\hbar$ on Spatio-Temporal Graph and Local Attention. Fig. 14 provides a visual representation of the attention distribution for reference tokens under various hyperparameter ($\hbar$) settings. In Fig. 14, the model’s capacity to differentiate between ‘cat’ and ‘dog’ reference tokens is displayed across a range of $\hbar$ values. At lower $\hbar$ values, the attention is more dispersed, indicating a wider, albeit less precise, focus on the relevant subjects within the image. As $\hbar$ increases, the attention distribution becomes more focused, suggesting a refined concentration that enhances the model’s ability to accurately distinguish between the cat and dog within the scene. However, it’s important to note that if $\hbar$ is excessively large, the reference local token may not effectively capture crucial local information between adjacent frames. This could result in under utilization of the temporal information between video frames, potentially leading to a decrease in the model’s feature table modeling ability. These figures collectively highlight the delicate balance required when setting the value of $\hbar$ to optimize the model’s performance in identifying and focusing on specific subjects within complex visual scenes.
@@ 5 Conclusion
This paper introduces a novel spatio-temporal graph transformer (STGT) module for video-language alignment. By combining the strengths of both graph and transformer, the STGT effectively learns the spatial and temporal features of videos, utilizing the spatio-temporal contexts provided by the spatio-temporal graph. Additionally, we propose a novel cross-similarity alignment loss (CSAL) to explore the inherent self-similarity via evaluating the corresponding two video-video and text-text pairs, further promoting the accuracy of video-text alignment. Experimental results on challenging downstream tasks demonstrate that our method surpasses existing state-of-the-art methods, achieving impressive performance in video-text retrieval and video question answering tasks.